OpenStudy (anonymous):

square root of 3q + 2 = Square root of 5

OpenStudy (anonymous):

\[\sqrt{3q+2}=\sqrt{5}?\]

OpenStudy (anonymous):

is that what you are asking?

OpenStudy (anonymous):

yes... @BigTeach2345

OpenStudy (anonymous):

square both sides to get rid of the square roots and then it will be easier to solve

OpenStudy (anonymous):

\[\sqrt{3q+2}^2=\sqrt{5}^2\] \[3q+2=5\]

OpenStudy (anonymous):

example?

OpenStudy (anonymous):

when you square a square root they cancel each other out! Do you see that?